OpenStudy (anonymous):

SOMEONE PLEASE HELP URGENT!!! what is the average rate of change from x=3 to x=5 for the function g(x)=log(x+1)? Round your answer to the nearest hundredth.

OpenStudy (unklerhaukus):

The average rate of change is \[\frac{\Delta g}{\Delta x} = \frac{g(5)-g(3)}{5-3}\] Can you find \(g(5)\), \(g(3)\)?

OpenStudy (anonymous):

is the answer 0.088?

OpenStudy (unklerhaukus):

i don't know the answers, show working

OpenStudy (unklerhaukus):

\[g(5)=\log(5+1)=\]

OpenStudy (anonymous):

0.155

OpenStudy (unklerhaukus):

i don't know what that is supposed to be